OpenStudy (fairy733189):

Which statement best describes the graph of -x3 - x2 + 4x + 4? t starts down on the left and goes up on the right and intersects the x-axis at x = -2, 2, and 4. It starts down on the left and goes up on the right and intersects the x-axis at x = -2, -1, and 2. It starts up on the left and goes down on the right and intersects the x-axis at x = -2, -1, and 2. It starts up on the left and goes down on the right and intersects the x-axis at x = -2, 2, and 4.

OpenStudy (wcrmelissa2001):

firstly factorise -x3-x2+4x+4 to find the x intercepts. Let f(x) = -x3-x2+4x+4 By trial and error, f(-1)=-(-1)^3-2(-1)+4(-1)+4)=0 Hence, (x+1) is a factor of f(x) By long division/inspection/synthetic division (whichever you're more comfortable with but I'm gonna use synthetic), f(x)÷(x+1)

OpenStudy (wcrmelissa2001):